EDITORIAL

Terry Goodman

This semester Robert Kennedy (our Coordinating Editor) and I are teaching a History of Mathematics course via satellite. The course, which consists of eleven class sessions (nine of these are broadcast sessions), is being broadcast to eleven public school sites around the state of Missouri. The audience for the course consists primarily of middle school mathematics teachers. Dr. Kennedy is the primary instructor for the course and I provide information for the teachers as to how they might use some of the topics from the course with their students. The format (TV) is quite different for Bob and me and we are having to make a number of adjustments in our teaching “styles”.

In one of our sessions Bob will discuss with the audience the “Three Greatest Mathematicians of All Time.” In discussing this course with other mathematicians we have found that they are often curious as to our choices for the three greatest mathematicians. Obviously, this is not an issue that can be resolved objectively — there will be many different opinions. Before we share with our readers who we consider to be the three greatest mathematicians we would like to solicit your input concerning this question. Please contact us by mail or e–mail with your views of the three greatest mathematicians and your rationale for your choices. In a future issue of MJMS, we will publish the results of this rather informal survey. We believe this will be interesting and informative.

In the coming year, we will be initiating more changes in MJMS — new cover design, new size, and perhaps a fourth issue each year. The input we have received so far concerning these changes has been positive and we appreciate hearing from you. Please continue to communicate with us concerning these and other improvements in our journal.

At this time, we would like to say thank you to those individuals who serve as referees for MJMS. We deeply appreciate their contribution to this journal and recognize that their role is of vital importance.

List of Referees

Hadi Y. Alkahby, Dillard University Charles J. Allard, Polo R-VII High School Mohammad K. Azarian, University of Evansville Seung–Jin Bang, Seoul, Korea Gerald Bergum, South Dakota State University John E. Bernard, University of Texas–Pan American Jim Bruening, Southeast Missouri State University Joseph Chance, University of Texas–Pan American Hang Chen, Central Missouri State University Bryan Dawson, Emporia State University Richard Delaware, University of Missouri–Kansas City Joseph B. Dence, University of Missouri–St. Louis Albert Dixon, College of the Ozarks John J. Edgell, Jr., Southwest Texas State University Vince Edmondson, Central Missouri State University Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Martin Erickson, Northeast Missouri State University Russell Euler, Northwest Missouri State University Joe Flowers, Northeast Missouri State University Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, Northwest Missouri State University Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh Dennis R. Harmon, Missouri Southern State College Stan Hartzler, Northwest Missouri State University Troy Hicks, University of Missouri–Rolla Yue–Chan Phoebe Ho, Central Missouri State University Larry K. Johnson, Central Missouri State University

List of Referees (cont.)

N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State College Roger Marty, Cleveland State University Wayne McDaniel, University of Missouri–St. Louis Gary McDonald, Northwest Missouri State University Rhonda L. McKee, Central Missouri State University Kandasamy Muthvel, University of Wisconsin–Oshkosh Ollie Nanyes, Bradley University Krzysztof Ostaszewski, University of Louisville Miguel Paredes, University of Texas–Pan American Bob Prielipp, University of Wisconsin–Oshkosh Don Redmond, Southern Illinois University Jeffrey A. Rosoff, Gustavus Adolphus College Prasanna Sahoo, University of Louisville Jau–Shyong Shiue, University of Nevada–Las Vagas Donald P. Skow, University of Texas–Pan American Jinhua Tao, Central Missouri State University Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Alvin Tinsley, Central Missouri State University Bobby Neal Winters, Pittsburg State University æ